OpenStudy (anonymous):

The product of a number decreased by 1 and increased by 8 is 22. Find all the numbers that satisfy this condition.

OpenStudy (anonymous):

Answer would be a solution set.

OpenStudy (anonymous):

lets call the number \(x\) so if you decrease it by 1 you get \(x-1\) and if you increase it by 8 you get \(x+8\) the product of the two numbers is \((x-1)(x+8)\) which you know is 22 so set \[(x-1)(x+8)=22\] and solve the resulting quadratic equation

OpenStudy (anonymous):

\[(x-1)(x+8)=22\\ \] get a quadratic

OpenStudy (anonymous):

multiply out, then set the whole thing equal to zero and solve

OpenStudy (anonymous):

using quadratic formula i got x= 3,-10

OpenStudy (anonymous):

yep.. two answers

OpenStudy (anonymous):

easy to check \((3-1)(3+8)=2\times 11=22\) and \((-10-1)(-10+8)=-11\times -2=22\)

OpenStudy (anonymous):

Thank you.